OpenStudy (anonymous):

Find how many 3 digit numbers are even? How many 3 digit numbers are composed of odd digits ?

OpenStudy (jt950):

the number can be repeated?

OpenStudy (dumbcow):

3 digit numbers range from 100 to 999 ..... that is 1000 total, half are even half are odd thus 500 3-digit num are even

OpenStudy (dumbcow):

there are 5 odd digits ....so each digit has 50% chance of being odd probability all 3 digits are odd = 1/2 * 1/2 *1/2 = 1/8 1000 * 1/8 = 125 125 3-digit num have all odd digits

OpenStudy (anonymous):

last digit = 5 options ( 0,2,4,6,8) , the rest 9 ( all no's) = 9*9*5 = 405

OpenStudy (anonymous):

odd digits = 5*5*5 = 125